OpenStudy (anonymous):

(5x-10) (2x-5) slove for x

OpenStudy (aum):

Is it equal to zero?

OpenStudy (anonymous):

5x-10=0 5x=10 x=2 2x-5+0 2x=5 x=5/2

OpenStudy (anonymous):

* 2x-5=0

OpenStudy (anonymous):

thanks so much(;

OpenStudy (anonymous):

mhm :)